i'm stuck solving for "D" in this problem.
c=54*{D^-(1/3)}*b
thank you for any and all help

Answer by stanbon(75887)

c=54*{D^-(1/3)}*b
------
(D^(-1/3)) = c/(54b)
----
D = [c/54b]^(-3)
=====
D = (157464b^3)/c^3
------
D = 157464(b/c)^3
==========================
Cheers,
Stan H.
===============